OpenStudy (anonymous):

Determine the nature of the solution of equations. x^2+19=0

OpenStudy (anonymous):

the choices are a. 2 imaginary solutions b. 1 real solution c. 2 real solutions

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[\left\{x\to -i \sqrt{19},x\to i \sqrt{19}\right\} \]

OpenStudy (anonymous):

np